Hindelang Inc is considering a project that has the following cash fow and WACC data What is the projects MIRR? Note that a projects projected MIRR can be less han the tWACC (and even negative) ), in whuch case it will be rejected WACC: 12.50% Year $360 Cash flows $850 $300 $320 $340 015.00% О 12.83% 01667% 1333% 2064% Previous Next Quiz saved at 8:24am·Submit Quiz DOLL Q WE Caps Lock A S D
Answer #1

Answer: MIRR : 16.67 %

Present value of costs ( PVC ) =   850.

Terminal Value ( TV ) =300 x ( 1.125) 3 + 320 ( 1.125) 2 + 340 x ( 1.1250 ) + 360 = 427.15 + 405 + 382.50 + 360 = 1,574.65

MIRR is obtained as follows:

PVC = TV / ( 1 + MIRR ) 4

or ( 1 + MIRR ) 4 = TV / PVC = 1,574.65 / 850 = 1.85253

MIRR = 0.1667 or 16.67 %